Assuming that a firm produces 200,000 vehicles when the market price is set at $ 30,000 per unit in a perfectly competitive market. The total cost for producing the 200,000 vehicles is $5,000,000.

  1. What is the firm’s total profit when it produces all the 200,000 vehicles?
  2. What level of production (number of vehicles produced) will bring the profit to $0?
  3. Should the firm stop producing when the total profit reaches $0? Explain based on the theory of ‘profit optimization’!
  4. What happens to the firm’s profit level when its production is limited to 100,000 vehicles?
  5. What should be the value of the optimizing Marginal Cost considering that the market price is set at $ 30,000?
1
Expert's answer
2021-04-13T07:15:26-0400

1. The firm’s total profit when it produces all the 200,000 vehicles is:

TP = 30,000×200,000 - 5,000,000 = $995,000,000,000.

2. The level of production (number of vehicles produced) that will bring the profit to $0 is: TC/P = 5,000,000/30,000 = 166.67 units.

3. The firm should not stop producing when the total profit reaches $0, because all firms earn zero profits in the long run.

4. The firm’s profit level when its production is limited to 100,000 vehicles will decrease.

5. The value of the optimizing Marginal Cost considering that the market price is set at $ 30,000 should be: MC = P = $30,000.